OpenStudy (anonymous):

Evaluate x/y for x=2/3 and y=6/7

OpenStudy (anonymous):

i dont understand how to do this... someone pls help

OpenStudy (bibby):

division = multiplication by the reciprocal

OpenStudy (anonymous):

So to divide them you multiply them by the recip. 6/7 * 3/2

OpenStudy (anonymous):

so i would cross multiply sort of?

OpenStudy (anonymous):

\[\frac {6}{7} * \frac {3} {2} = \frac {6/7} {2/3}\]

OpenStudy (anonymous):

okay so the answer would be 18/14?

OpenStudy (anonymous):

or 7/9?

OpenStudy (anonymous):

\[\frac {a}{b} * \frac {c}{d} = \frac {a*c} {b*d}\]

OpenStudy (anonymous):

Do they want it simplified?

OpenStudy (anonymous):

18/14 would simplify to what?

OpenStudy (anonymous):

9/7

OpenStudy (anonymous):

is that right?

OpenStudy (anonymous):

Exactly!